# TimeSiam: A Pre-Training Framework for Siamese Time-Series Modeling

### v1

@@ 4.6 Analysis Experiment
As an important finetuning setting, we also experiment with the linear probing, where we fix the pre-trained encoder and only finetune the newly added projector at the end of the model. Figure 5 shows that TimeSiam outperforms other baselines on linear probing performance.
Figure 5: Linear probing on in-domain forecasting setting. Average results (MSE) are reported. Full results are in Table 16.
@@ A.3 Model Configuration
In addition, we performed a hyperparameter search for all baselines, adhering to their official configuration in the in-domain setting. For Siamese encoders, we explored various configurations by adjusting the number of encoder layers ($e_{\text{layers}}$) and decoder layers ($d_{\text{layers}}$) from $\{1,2,3,4\}$, selecting hidden dimensions ($D$) from $\{16,32,64,128,256,512\}$ and attention heads ($n_{\text{heads}}$) from $\{8,16,32\}$. In the case of TCN models, we investigated different numbers of residual blocks, considering configurations of $\{5,8,10\}$.
@@ B.2 Classification Datasets
(1) AD has electroencephalography (EEG) recordings from 12 Alzheimer’s patients and 11 healthy controls. Each patient has around 30 trials, each lasting for 5 seconds with 1280 timestamps (sampled at 256Hz) and includes 16 channels.
(2) PTB has electrocardiogram (ECG) recordings from 290 patients with 15 channels sampled at 1000 Hz. This paper focuses on a subset of the dataset that includes 198 patients with heart diseases: Myocardial infarction and healthy controls.
(3) TDBrain monitors brain signals of 1274 patients with 33 channels during EC (Eye closed) and EO (Eye open) tasks. It includes 60 types of diseases, but this paper focuses on a subset of 25 Parkinson’s disease patients and 25 healthy controls. Only the EC task trials are used for representation learning.
@@ Appendix E Multiple Lineages Representation Visualization
We employ Principal Components Analysis (PCA) to elucidate the distribution of temporal representations on the ECL dataset. It is worth noting that when time series is fed into a pre-trained Siamese network with different lineage embeddings, the model generates divergent temporal representations that representations derived from the same lineage embeddings tend to be closely clustered together, while representations from different lineage embeddings exhibit significant dissimilarity. This observation effectively validates the efficacy of combining a pre-trained Siamese network with lineage embeddings in enhancing the diversity of representations.
@@ H.2 Different Datasets
Methods Random init. TS2Vec CoST LaST TFC TST Ti-MAE SimMTM TimeSiame
ETTh1 96 0.420 0.423 0.381 0.400 0.383 0.405 0.396 0.413 0.399 0.420 0.377 0.401 0.396 0.415 0.367 0.389 0.378 0.401
192 0.465 0.449 0.421 0.427 0.434 0.437 0.457 0.451 0.444 0.449 0.432 0.436 0.440 0.443 0.424 0.423 0.422 0.430
336 0.504 0.470 0.468 0.452 0.474 0.460 0.507 0.478 0.479 0.467 0.475 0.461 0.481 0.462 0.473 0.456 0.459 0.452
720 0.502 0.492 0.553 0.507 0.535 0.509 0.516 0.508 0.491 0.490 0.525 0.500 0.475 0.481 0.494 0.493 0.459 0.466
Avg 0.473 0.458 0.456 0.447 0.457 0.453 0.469 0.463 0.453 0.457 0.452 0.450 0.448 0.450 0.440 0.440 0.429 0.437
ETTh2 96 0.297 0.345 0.297 0.343 0.288 0.342 0.294 0.345 0.302 0.345 0.304 0.358 0.300 0.353 0.299 0.352 0.293 0.345
192 0.388 0.400 0.366 0.392 0.367 0.392 0.379 0.395 0.369 0.392 0.379 0.403 0.372 0.396 0.380 0.398 0.370 0.392
336 0.426 0.434 0.416 0.430 0.413 0.429 0.423 0.436 0.412 0.428 0.412 0.432 0.418 0.431 0.422 0.432 0.410 0.424
720 0.431 0.446 0.424 0.447 0.428 0.446 0.445 0.460 0.428 0.446 0.438 0.457 0.427 0.447 0.428 0.449 0.418 0.440
Avg 0.385 0.406 0.376 0.403 0.374 0.402 0.385 0.409 0.378 0.403 0.383 0.413 0.379 0.407 0.382 0.408 0.373 0.400
ETTm1 96 0.330 0.368 0.325 0.364 0.348 0.377 0.345 0.381 0.353 0.378 0.319 0.360 0.325 0.363 0.317 0.356 0.319 0.360
192 0.369 0.385 0.370 0.389 0.367 0.387 0.372 0.391 0.361 0.384 0.360 0.387 0.363 0.385 0.362 0.387 0.353 0.379
336 0.400 0.407 0.405 0.415 0.404 0.414 0.412 0.420 0.392 0.406 0.391 0.408 0.396 0.409 0.387 0.405 0.383 0.402
720 0.460 0.439 0.471 0.452 0.460 0.447 0.462 0.448 0.448 0.440 0.449 0.445 0.452 0.438 0.443 0.438 0.440 0.436
Avg 0.390 0.400 0.393 0.405 0.395 0.406 0.398 0.410 0.389 0.402 0.380 0.400 0.384 0.399 0.377 0.397 0.374 0.394
ETTm2 96 0.175 0.258 0.174 0.415 0.181 0.269 0.177 0.258 0.281 0.327 0.181 0.265 0.175 0.261 0.175 0.262 0.175 0.261
192 0.247 0.307 0.247 0.306 0.247 0.312 0.252 0.309 0.241 0.302 0.247 0.309 0.241 0.303 0.244 0.307 0.241 0.303
336 0.309 0.345 0.306 0.345 0.309 0.348 0.307 0.344 0.304 0.343 0.314 0.354 0.301 0.341 0.312 0.351 0.300 0.341
720 0.408 0.403 0.427 0.415 0.408 0.406 0.404 0.402 0.404 0.403 0.408 0.407 0.398 0.397 0.410 0.408 0.399 0.398
Avg 0.285 0.328 0.289 0.370 0.286 0.334 0.285 0.328 0.281 0.327 0.288 0.334 0.279 0.326 0.285 0.332 0.279 0.326
Weather 96 0.177 0.218 0.174 0.216 0.171 0.214 0.170 0.212 0.177 0.218 0.177 0.221 0.175 0.218 0.184 0.220 0.171 0.213
192 0.225 0.259 0.220 0.257 0.218 0.255 0.215 0.253 0.222 0.257 0.223 0.260 0.222 0.256 0.217 0.255 0.217 0.253
336 0.278 0.297 0.276 0.297 0.273 0.295 0.272 0.295 0.277 0.296 0.279 0.301 0.278 0.299 0.273 0.296 0.272 0.293
720 0.354 0.348 0.352 0.346 0.350 0.344 0.349 0.344 0.353 0.346 0.355 0.350 0.353 0.346 0.348 0.344 0.348 0.343
Avg 0.259 0.281 0.256 0.279 0.253 0.277 0.252 0.276 0.257 0.279 0.259 0.283 0.257 0.280 0.256 0.279 0.252 0.276
Exchange 96 0.084 0.201 0.084 0.201 0.090 0.208 0.096 0.220 0.083 0.201 0.098 0.218 0.083 0.200 0.083 0.202 0.084 0.203
192 0.187 0.307 0.185 0.306 0.179 0.301 0.190 0.313 0.173 0.296 0.187 0.308 0.186 0.307 0.182 0.303 0.176 0.300
336 0.337 0.422 0.328 0.415 0.332 0.416 0.409 0.455 0.332 0.418 0.330 0.418 0.327 0.415 0.346 0.427 0.310 0.404
720 0.858 0.695 0.856 0.696 0.854 0.698 1.035 0.749 0.860 0.698 0.925 0.731 0.882 0.708 0.831 0.689 0.842 0.690
Avg 0.367 0.406 0.363 0.405 0.364 0.406 0.433 0.434 0.362 0.403 0.385 0.419 0.370 0.408 0.361 0.405 0.353 0.399
ECL 96 0.193 0.291 0.175 0.268 0.178 0.269 0.183 0.275 0.171 0.263 0.171 0.267 0.181 0.271 0.164 0.255 0.164 0.245
192 0.199 0.297 0.183 0.275 0.185 0.275 0.190 0.281 0.188 0.277 0.181 0.276 0.197 0.277 0.178 0.268 0.173 0.256
336 0.216 0.312 0.199 0.292 0.202 0.292 0.205 0.296 0.205 0.291 0.197 0.291 0.200 0.293 0.190 0.280 0.189 0.275
720 0.257 0.345 0.240 0.324 0.245 0.326 0.248 0.330 0.244 0.322 0.237 0.325 0.205 0.326 0.235 0.318 0.229 0.310
Avg 0.216 0.311 0.199 0.290 0.203 0.291 0.207 0.296 0.202 0.288 0.197 0.290 0.196 0.292 0.192 0.280 0.189 0.272
Traffic 96 0.472 0.305 0.449 0.291 0.458 0.294 0.506 0.330 0.465 0.301 0.478 0.292 0.463 0.295 0.442 0.285 0.429 0.279
192 0.474 0.304 0.457 0.293 0.465 0.297 0.503 0.326 0.470 0.311 0.469 0.316 0.470 0.299 0.452 0.305 0.442 0.282
336 0.491 0.331 0.474 0.301 0.480 0.304 0.517 0.332 0.498 0.320 0.482 0.323 0.486 0.321 0.473 0.322 0.456 0.288
720 0.523 0.327 0.509 0.319 0.515 0.321 0.552 0.349 0.514 0.326 0.516 0.327 0.504 0.337 0.497 0.331 0.486 0.307
Avg 0.490 0.317 0.472 0.301 0.480 0.304 0.520 0.334 0.487 0.315 0.486 0.315 0.481 0.313 0.466 0.311 0.453 0.289

### v2

@@ 4.6 Analysis Experiment
As an important finetuning setting, we also experiment with the linear probing, where we fix the pre-trained encoder and only finetune the newly added projector at the end of the model. Figure 5 illustrates that TimeSiam demonstrates superior performance compared to other baselines in terms of overall linear probing performance. Interestingly, by only fine-tuning the model head, the average forecasting performance across the four ETT subsets is already comparable with the results obtained through full fine-tuning, and significantly outperforms training from random initialization (MSE: 0.365 vs. 0.383). This finding further validates the effectiveness of TimeSiam in learning generalizable representations for various downstream tasks.
Figure 5: Linear probing on in-domain forecasting setting. Average results (MSE) are reported. Full results are shown in Table 17.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Key Research and Development Plan (2021YFB1715200), the National Natural Science Foundation of China (62022050 and U2342217), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software. We are grateful to our colleagues Yipeng Huang and Zhiyao Cen for their support in the experimental efficiency.
Impact Statement
This paper focus on developing practical time series pre-training methods. We presents a novel approach based on Siamese networks, which could provide some inspiration for future research. The experimental results demonstrate the effectiveness of our approach across various domains and its potential value for real-world applications. It is essential to note that our work focuses solely on scientific issues, and we also ensure that ethical considerations are carefully taken into account. All the medical-related datasets are publicly available for scientific research. Thus, we believe that there is no ethical risk associated with our research.
@@ A.3 Model Configuration
In addition, we performed a hyperparameter search for all baselines, adhering to their official configuration in the in-domain setting. For Siamese encoders, we explored various configurations by adjusting the number of encoder layers ($e_{\text{layers}}$) and decoder layers ($d_{\text{layers}}$) from $\{1,2,3,4\}$, selecting hidden dimensions ($d_{\text{model}}$) from $\{16,32,64,128,256,512\}$ and attention heads ($n_{\text{heads}}$) from $\{8,16,32\}$. In the case of TCN models, we investigated different numbers of residual blocks, considering configurations of $\{5,8,10\}$. During the fine-tuning stage, we carefully consider the learning rate (lr) from $\{1\rm{e}\mbox{-}3,5\rm{e}\mbox{-}4,1\rm{e}\mbox{-}4,1\rm{e}\mbox{-}5\}$, and head dropout (dropout) from $\{0,0.1,0.2,0.3\}$ in order to enhance the adaptability of our pretrained model to diverse datasets.
@@ B.2 Classification Datasets
(1) AD  (Escudero et al., 2006) has electroencephalography (EEG) recordings from 12 Alzheimer’s patients and 11 healthy controls. Each patient has around 30 trials, each lasting for 5 seconds with 1280 timestamps (sampled at 256Hz) and includes 16 channels.
(2) PTB  (Goldberger et al., 2000) has electrocardiogram (ECG) recordings from 290 patients with 15 channels sampled at 1000 Hz. This paper focuses on a subset of the dataset that includes 198 patients with heart diseases: Myocardial infarction and healthy controls.
(3) TDBrain  (Van Dijk et al., 2022) monitors brain signals of 1274 patients with 33 channels during EC (Eye closed) and EO (Eye open) tasks. It includes 60 types of diseases, but this paper focuses on a subset of 25 Parkinson’s disease patients and 25 healthy controls. Only the EC task trials are used for representation learning.
@@ Appendix E Multiple Lineages Representation Visualization
We employ Principal Components Analysis (PCA) to elucidate the distribution of temporal representations on the ECL dataset. We will only train the learned lineage embeddings during the pre-training phase. However, during downstream fine-tuning or linear probing, we will keep them fixed and not update them. It is worth noting that the embedded feature will be the same without different lineage embeddings. However, when time series is fed into a pre-trained Siamese network with different lineage embeddings, the model generates divergent temporal representations that representations derived from the same lineage embeddings tend to be closely clustered together, while representations from different lineage embeddings exhibit significant dissimilarity. Upon visual analysis, we have observed that the representations generated based on the same data but with different lineage embeddings exhibit a high level of diversity. This observation effectively validates the effectiveness of combining a pre-trained Siamese network with different lineage embeddings, which can enlarge the representation diversity.
@@ H.2 Different Datasets
Methods Random init. CPC TNC TS2Vec CoST LaST TFC TST Ti-MAE SimMTM TimeSiame
Metric mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae mse mae
ETTh1 96 0.420 0.423 0.380 0.401 0.377 0.397 0.381 0.400 0.383 0.405 0.396 0.413 0.399 0.420 0.377 0.401 0.396 0.415 0.367 0.389 0.378 0.401
192 0.465 0.449 0.426 0.429 0.423 0.427 0.421 0.427 0.434 0.437 0.457 0.451 0.444 0.449 0.432 0.436 0.440 0.443 0.424 0.423 0.422 0.430
336 0.504 0.470 0.465 0.451 0.471 0.453 0.468 0.452 0.474 0.460 0.507 0.478 0.479 0.467 0.475 0.461 0.481 0.462 0.473 0.456 0.459 0.452
720 0.502 0.492 0.488 0.479 0.508 0.485 0.553 0.507 0.535 0.509 0.516 0.508 0.491 0.490 0.525 0.500 0.475 0.481 0.494 0.493 0.459 0.466
Avg 0.473 0.458 0.440 0.440 0.445 0.441 0.456 0.447 0.457 0.453 0.469 0.463 0.453 0.457 0.452 0.450 0.448 0.450 0.440 0.440 0.429 0.437
ETTh2 96 0.297 0.345 0.313 0.364 0.291 0.343 0.297 0.343 0.288 0.342 0.294 0.345 0.302 0.345 0.304 0.358 0.300 0.353 0.299 0.352 0.293 0.345
192 0.388 0.400 0.392 0.412 0.366 0.393 0.366 0.392 0.367 0.392 0.379 0.395 0.369 0.392 0.379 0.403 0.372 0.396 0.380 0.398 0.370 0.392
336 0.426 0.434 0.438 0.449 0.427 0.438 0.416 0.430 0.413 0.429 0.423 0.436 0.412 0.428 0.412 0.432 0.418 0.431 0.422 0.432 0.410 0.424
720 0.431 0.446 0.460 0.470 0.433 0.451 0.424 0.447 0.428 0.446 0.445 0.460 0.428 0.446 0.438 0.457 0.427 0.447 0.428 0.449 0.418 0.440
Avg 0.385 0.406 0.401 0.424 0.379 0.406 0.376 0.403 0.374 0.402 0.385 0.409 0.378 0.403 0.383 0.413 0.379 0.407 0.382 0.408 0.373 0.400
ETTm1 96 0.330 0.368 0.324 0.361 0.323 0.361 0.325 0.364 0.348 0.377 0.345 0.381 0.353 0.378 0.319 0.360 0.325 0.363 0.317 0.356 0.319 0.360
192 0.369 0.385 0.368 0.382 0.366 0.383 0.370 0.389 0.367 0.387 0.372 0.391 0.361 0.384 0.360 0.387 0.363 0.385 0.362 0.387 0.353 0.379
336 0.400 0.407 0.403 0.405 0.399 0.405 0.405 0.415 0.404 0.414 0.412 0.420 0.392 0.406 0.391 0.408 0.396 0.409 0.387 0.405 0.383 0.402
720 0.460 0.439 0.461 0.439 0.457 0.439 0.471 0.452 0.460 0.447 0.462 0.448 0.448 0.440 0.449 0.445 0.452 0.438 0.443 0.438 0.440 0.436
Avg 0.390 0.400 0.389 0.397 0.386 0.397 0.393 0.405 0.395 0.406 0.398 0.410 0.389 0.402 0.380 0.400 0.384 0.399 0.377 0.397 0.374 0.394
ETTm2 96 0.175 0.258 0.196 0.281 0.187 0.261 0.174 0.261 0.181 0.269 0.177 0.258 0.281 0.327 0.181 0.265 0.175 0.261 0.175 0.262 0.175 0.261
192 0.247 0.307 0.261 0.323 0.241 0.302 0.247 0.306 0.247 0.312 0.252 0.309 0.241 0.302 0.247 0.309 0.241 0.303 0.244 0.307 0.241 0.303
336 0.309 0.345 0.302 0.343 0.313 0.363 0.306 0.345 0.309 0.348 0.307 0.344 0.304 0.343 0.314 0.354 0.301 0.341 0.312 0.351 0.300 0.341
720 0.408 0.403 0.399 0.397 0.408 0.407 0.427 0.415 0.408 0.406 0.404 0.402 0.404 0.403 0.408 0.407 0.398 0.397 0.410 0.408 0.399 0.398
Avg 0.285 0.328 0.290 0.336 0.287 0.333 0.289 0.332 0.286 0.334 0.285 0.328 0.281 0.327 0.288 0.334 0.279 0.326 0.285 0.332 0.279 0.326
Weather 96 0.177 0.218 0.193 0.230 0.191 0.231 0.174 0.216 0.171 0.214 0.170 0.212 0.177 0.218 0.177 0.221 0.175 0.218 0.184 0.220 0.171 0.213
192 0.225 0.259 0.238 0.267 0.237 0.267 0.220 0.257 0.218 0.255 0.215 0.253 0.222 0.257 0.223 0.260 0.222 0.256 0.217 0.255 0.217 0.253
336 0.278 0.297 0.292 0.306 0.292 0.305 0.276 0.297 0.273 0.295 0.272 0.295 0.277 0.296 0.279 0.301 0.278 0.299 0.273 0.296 0.272 0.293
720 0.354 0.348 0.364 0.352 0.361 0.349 0.352 0.346 0.350 0.344 0.349 0.344 0.353 0.346 0.355 0.350 0.353 0.346 0.348 0.344 0.348 0.343
Avg 0.259 0.281 0.272 0.289 0.270 0.288 0.256 0.279 0.253 0.277 0.252 0.276 0.257 0.279 0.259 0.283 0.257 0.280 0.256 0.279 0.252 0.276
Exchange 96 0.084 0.201 0.085 0.197 0.086 0.203 0.084 0.201 0.090 0.208 0.096 0.220 0.083 0.201 0.098 0.218 0.083 0.200 0.083 0.202 0.084 0.203
192 0.187 0.307 0.187 0.308 0.180 0.301 0.185 0.306 0.179 0.301 0.190 0.313 0.173 0.296 0.187 0.308 0.186 0.307 0.182 0.303 0.176 0.300
336 0.337 0.422 0.332 0.422 0.329 0.416 0.328 0.415 0.332 0.416 0.409 0.455 0.332 0.418 0.330 0.418 0.327 0.415 0.346 0.427 0.310 0.404
720 0.858 0.695 0.867 0.694 0.851 0.694 0.856 0.696 0.854 0.698 1.035 0.749 0.860 0.698 0.925 0.731 0.882 0.708 0.831 0.689 0.842 0.690
Avg 0.367 0.406 0.368 0.405 0.362 0.404 0.363 0.405 0.364 0.406 0.433 0.434 0.362 0.403 0.385 0.419 0.370 0.408 0.361 0.405 0.353 0.399
ECL 96 0.193 0.291 0.190 0.287 0.190 0.277 0.175 0.268 0.178 0.269 0.183 0.275 0.171 0.263 0.171 0.267 0.181 0.271 0.164 0.255 0.164 0.245
192 0.199 0.297 0.204 0.290 0.194 0.283 0.183 0.275 0.185 0.275 0.190 0.281 0.188 0.277 0.181 0.276 0.197 0.277 0.178 0.268 0.173 0.256
336 0.216 0.312 0.227 0.300 0.211 0.299 0.199 0.292 0.202 0.292 0.205 0.296 0.205 0.291 0.197 0.291 0.200 0.293 0.190 0.280 0.189 0.275
720 0.257 0.345 0.257 0.347 0.254 0.334 0.240 0.324 0.245 0.326 0.248 0.330 0.244 0.322 0.237 0.325 0.205 0.326 0.235 0.318 0.229 0.310
Avg 0.216 0.311 0.220 0.306 0.212 0.298 0.199 0.290 0.203 0.291 0.207 0.296 0.202 0.288 0.197 0.290 0.196 0.292 0.192 0.280 0.189 0.272
Traffic 96 0.472 0.305 0.449 0.487 0.483 0.309 0.309 0.291 0.458 0.294 0.506 0.330 0.465 0.301 0.478 0.292 0.463 0.295 0.442 0.285 0.429 0.279
192 0.474 0.304 0.505 0.315 0.495 0.311 0.457 0.293 0.465 0.297 0.503 0.326 0.470 0.311 0.469 0.316 0.470 0.299 0.452 0.305 0.442 0.282
336 0.491 0.331 0.514 0.343 0.504 0.333 0.474 0.301 0.480 0.304 0.517 0.332 0.498 0.320 0.482 0.323 0.486 0.321 0.473 0.322 0.456 0.288
720 0.523 0.327 0.511 0.351 0.521 0.341 0.509 0.319 0.515 0.321 0.552 0.349 0.514 0.326 0.516 0.327 0.504 0.337 0.497 0.331 0.486 0.307
Avg 0.490 0.317 0.504 0.330 0.501 0.324 0.472 0.301 0.480 0.304 0.520 0.334 0.487 0.315 0.486 0.315 0.481 0.313 0.466 0.311 0.453 0.289
@@
CPC $\text{77.40}_{\pm 7.28}$ $\text{79.91}_{\pm 4.35}$ $\text{78.52}_{\pm 6.18}$ $\text{77.09}_{\pm 7.65}$ $\text{89.81}_{\pm 2.98}$ $\text{89.49}_{\pm 3.20}$
TNC $\text{78.58}_{\pm 6.21}$ $\text{81.10}_{\pm 4.09}$ $\text{79.97}_{\pm 5.50}$ $\text{78.43}_{\pm 6.35}$ $\text{92.26}_{\pm 2.38}$ $\text{92.10}_{\pm 2.60}$
@@
CPC $\text{85.19}_{\pm 2.99}$ $\text{85.35}_{\pm 2.88}$ $\text{85.19}_{\pm 2.99}$ $\text{85.17}_{\pm 3.01}$ $\text{93.50}_{\pm 2.55}$ $\text{93.68}_{\pm 2.50}$
TNC $\text{85.21}_{\pm 1.92}$ $\text{86.49}_{\pm 1.86}$ $\text{85.21}_{\pm 1.92}$ $\text{85.08}_{\pm 1.95}$ $\text{95.77}_{\pm 1.30}$ $\text{95.95}_{\pm 1.25}$
@@
CPC $\text{88.30}_{\pm 3.07}$ $\text{88.90}_{\pm 1.00}$ $\text{81.54}_{\pm 6.51}$ $\text{83.75}_{\pm 5.67}$ $\text{89.86}_{\pm 3.87}$ $\text{88.68}_{\pm 2.89}$
TNC $\text{90.53}_{\pm 2.92}$ $\text{89.01}_{\pm 2.87}$ $\text{87.06}_{\pm 5.22}$ $\text{87.82}_{\pm 4.13}$ $\text{93.12}_{\pm 2.21}$ $\text{91.01}_{\pm 1.55}$